# Multimodal Large Language Models to Support Real-World Fact-Checking

### v1

@@ 4.5 RQ5: Calibration Evaluation
Calibration refers to the alignment between a model’s predicted probabilities (or confidence levels) and the actual outcomes. It reflects the model’s knowledge-awareness and ability to accurately estimate the confidence in its responses Geng et al. (2023). In this evaluation, we use verbalized confidence, which is readily accessible. Figure 4 displays the confidence distributions alongside calibration curves for GPT-4V and LLaVA(13b), breaking down the data into 10% confidence intervals. For instance, a value of 95 corresponds to confidence levels within the bracket. GPT-4V’s confidence scores are largely clustered in the interval, whereas LLaVA(13b)’s scores are more frequently found in the range. The calibration curves reveal that both models exhibit a positive correlation between confidence levels and accuracy, with GPT-4V’s calibration closely mirroring the ideal curve (illustrated by the dashed line), indicating well-calibrated confidence scores. Conversely, LLaVA(13b)’s curve suggests a propensity for overconfidence. The Expected Calibration Errors (ECEs, Guo et al. 2017), detailed in Table 6, further substantiate that GPT-4V maintains superior calibration.
ECE |
Fauxtography |
COSMOS |
MOCHEG |
Post-4V |
GPT-4V |
7.76 |
4.36 |
4.88 |
11.83 |
LLaVA |
30.20 |
15.26 |
22.82 |
28.45 |
4.6 RQ6: Taxonomy of Failure Reasons
Upon examining GPT-4V’s explanations, we identify six primary categories of failure reasons. Failures attributed to Lack of Knowledge occur when responses are based on trivial or incomplete information, often contradicting common sense. Recognition Errors arise from the model’s visual module misinterpreting images, including failing to identify altered or AI-generated visuals. Reasoning Errors occur when the model incorrectly interprets images due to factors like unnatural distortions or unusual perspectives, or there are self-contradictions in reasoning. Failures due to Partial Facts happen when the model addresses only a subset of the facts present in a textual claim. Image Issues involve failures related to poor image quality, such as blurriness or incompleteness, affecting the model’s ability to recognize content accurately. Ambiguity Claims involve semantic uncertainties, where a term could reference multiple entities. Examples beyond the lack of knowledge are presented in Figure 5. In subfigure (a), GPT-4V misinterprets scars from a dog’s burn as being painted or digitally altered. In subfigure (b), it correctly identifies the locations of Notre Dame and the Eiffel Tower but engages in erroneous reasoning. In subfigure (c), the model understands that a kiss can release hormones like oxytocin and knows that MR images cannot capture chemical reactions, with its predictions primarily influenced by the first fact. In subfigure (d), the model cannot discern that the scene depicted is related to posture training based on the image’s content. In subfigure (e), "Playboy" has multiple meanings, which could refer to a publisher or magazine’s name. We also manually tall the frequency of these errors across various datasets, as detailed in Table 7. Overall, the primary cause of errors is attributed to Lack of Knowledge, suggesting that augmenting the system with additional evidence could enhance its performance. There is room for improving the model’s ability to understand and reason about AI-generated or distorted images.
Know |
Recog |
Reas |
Partial |
Image |
Ambi |
| Fauxtography |
59.3 |
20.3 |
17.2 |
1.6 |
1.6 |
0.0 |
| COSMOS |
80.3 |
5.9 |
9.8 |
2.0 |
0 |
2.0 |
| MOCHEG |
73.2 |
3.6 |
7.1 |
3.6 |
5.4 |
7.1 |
| Post-4V |
85.2 |
11.1 |
0.0 |
0.0 |
0.0 |
3.7 |
We also explore the utilization of GPT-4 for detecting nonfactual content within explanations when fact-checking articles are available. We design a prompt comprising two steps: removing sentences containing no facts and predicting whether the reference article supports the sentence. Due to the limited content space of the article, we do not present the specific prompt. Figure 6 showcases two instances where, although the predictions are correct, the facts within the explanations are inaccurate. This underscores the necessity for further examination of their veracity. The photograph on the left, featuring a young girl named Fouzia, was captured in Morocco by Giacomo Pirozzi, a photographer for the United Nations Children’s Fund (UNICEF). However, GPT-4V incorrectly attributes it to photographer Terri Gold in Yemen, 1998. The right photo is incorrectly identified as from the movie "Platoon," actually, it is from the movie "Tropic Thunder." The returned sentences also contained substantial image descriptions not present in the original fact-checking articles. Improving the automatic assessment of the factuality in explanations is a focus for our future work.
@@ 5 Conclusion and Future Work
We investigated the capabilities of MLLMs to fact-check real-world claims, relying solely on their parametric knowledge and reasoning capabilities without external references. We proposed an evaluation framework, designed various experiments to address the research questions. Our results indicated that state-of-the-art MLLMs, such as GPT-4V, have the potential to assist professional fact-checkers: they can enhance their efficiency by providing reference predictions, valuable clues, and explanatory insights, together with confidence. Open-source models exhibit sensitivity to text, with different prompts revealing various biases. Moreover, they are susceptible to being deceived by altered images, and even in the absence of images, they tend to produce hallucinations pertaining to image descriptions. Our work demonstrates both the potential and limitations of MLLMs, laying the groundwork for further leveraging MLLMs to support fact-checking.
In future work, we plan to investigate how to enhance the fact-checking capabilities of models, focusing not only on improving accuracy but also on strengthening robustness and augmenting the knowledge reasoning and the verification abilities required for fact-checking. We further plan to explore how MLLMs can better support fact-checking when connected to external knowledge.

### v2

@@ 5.1 Comprehensive Results
The left part of Figure 3 displays the confidence distributions of GPT-4V and LLaVA(13b), breaking down the data into 10% confidence intervals. For instance, a value of 95 corresponds to confidence levels within the bracket. GPT-4V’s confidence scores are largely clustered in the interval, whereas LLaVA(13b)’s scores are more frequently found in the range. The right calibration curves reveal that both GPT-4V and LLaVA exhibit a positive correlation between confidence levels and accuracy, with GPT-4V’s calibration closely mirroring the ideal curve (illustrated by the dashed line), indicating well-calibrated confidence scores. Conversely, LLaVA(13b)’s curve suggests a propensity for overconfidence. We find that ICL does not result in better calibration of LLaVA’s verbalized confidence as their curves are very close.
5.2 Explanation-based Analysis
5.2.1 Case Studies
Figure 4 displays totally 4 cases with the original multimodal claims and corresponding responses . Rows labeled "GPT-4V w. Image," "LLaVA w. Image," and "LLaVA w. Image 1 example" correspond to the results of GPT-4V, LLaVA(13b), and LLaVA+ICL-1, respectively. It can be observed that GPT-4V’s explanation contains a richer set of facts. For instance, it explicitly states, "there was never an official campaign for Michelle Obama in the 2020 election," in the 1st case, and in the 4th case, it provides detailed background information about Margaret Hamilton, including her role in "leading the development of onboard flight software for the Apollo mission." Moreover, GPT-4V adeptly incorporates details from the images, such as referencing a "Google image search: Creepy Joe Biden" in the 3rd image.
However, LLaVA(13b) lacks the capability to counteract manipulated images, as evidenced in the 1st and 3rd cases. Additionally, LLaVA(13b)’s explanations are typically brief, often limited to a description of the image followed by a conclusion about its veracity. It shows a lack of effective reasoning, rendering the explanations less comprehensive than GPT-4V. We find ICL enhances LLaVA(13b)’s fact-checking capabilities. For instance, the model accurately analyzes the 3rd image, with "a genuine photograph of Joe Biden and an elderly woman, but the gun and menacing pose are added elements that are not part of the original scene." In the 4th case, LLaVA+ICL-1 provides a more detailed description of Margaret Hamilton’s background and states that "the image is an authentic representation of her contribution." However, it still fails in the 1st case with ICL.
5.2.2 Analysis of Explanation Length
For convenience, we refer to the number of sentences in the explanation part of a response as the explanation length. The average length of model explanations with different approaches is shown in Figure 5. We use NLTK to segment the sentences. GPT-4V provides the highest number of sentences in explanations across various datasets, with a minimum of 3.57 sentences in COSMOS and a maximum of 3.88 in Fauxtography. The value of LLaVA+PE represents the average length of explanations under different prompts. ICL prompts LLaVA to generate more sentences; however, providing two examples (LLaVA+ICL-2) does not yield longer explanations than providing just one (LLaVA+ICL-1), the average length even decreases.
We further analyze the relationship between the model’s prediction accuracy and the explanation length, as depicted in Figure 6, where we display explanation lengths ranging from 2 to 7, encompassing the most samples generated by various models. Typically, different models achieve peak accuracy when the explanation length is 5 or 6 sentences. Specifically, on COSMOS, the accuracy of the different models grows steadily as the explanation length increases from 2 to 5. Across the other three datasets, the left half of the curves all show varying degrees of ascent despite noticeable fluctuations. GPT-4V achieves its highest accuracy at an explanation length of 5 sentences on them.
| Fauxtography | COSMOS | MOCHEG | Post-4V | |||||||||||||||||||||
| T | F | U | O | T&F | All | T | F | U | O | T&F | All | T | F | U | O | T&F | All | T | F | U | O | T&F | All | |
| GPT-4V | 105 | 202 | 32 | 61 | 80.1 | 71.0 | 111 | 253 | 50 | 66 | 64.0 | 57.5 | 153 | 265 | 38 | 48 | 73.9 | 68.9 | 35 | 68 | 10 | 73 | 74.8 | 67.7 |
| LLaVA(7b) | 398 | 0 | 2 | 0 | 50.3 | 50.0 | 474 | 0 | 4 | 2 | 50.2 | 49.8 | 502 | 0 | 2 | 0 | 53.0 | 52.8 | 181 | 0 | 4 | 1 | 43.1 | 43.0 |
| LLaVA(13b) | 167 | 214 | 17 | 2 | 62.5 | 62.0 | 263 | 195 | 20 | 2 | 76.6 | 75.8 | 168 | 316 | 20 | 0 | 62.0 | 61.5 | 40 | 124 | 22 | 0 | 60.4 | 60.8 |
5.2.3 Reasons for Predicting False
To better understand the model’s fact-checking process, particularly why it labels multimodal claims as false, we explore two types of cases: checked cases, the model confirms that statements have been verified by organizations; manipulated cases, it indicates that images have been altered.
Checked Cases
We count the occurrences of phrases such as "been checked" and "been debunked" across different settings, with the results displayed in left graph in Figure 7. GPT-4V mentions these keywords most frequently across three public datasets, significantly more than others. Additionally, ICL prompts LLaVA to mention them more often. Notably, our ICL examples do not include checked cases. However, on Post-4V, several claims are considered fact-checked by LLaVA, but not by GPT-4V. The accuracy of GPT-4V and LLaVA+ICL on checked cases is and , respectively. The average True&False Accuracy for those that are not checked cases is only 84.4% and 65.7%, respectively.
Manipulated Cases
We notice that only the examples in the 3rd demonstration do not include manipulated cases. Therefore, we compare the impact of demonstrations on the number of manipulated cases, as shown in the right of Figure 7. Despite not accessing manipulated cases as example, the number significantly rises in demo3-1-exp1 and demo3-exp2, although these figures remain lower compared to other demonstrations. Thus, reasoning from the perspective of image manipulation is also not an explicit learning process. The accuracy of LLaVA is 67.6% for manipulated cases and 65.4% for non-manipulated cases, respectively. This demonstrates LLaVA’s potential to analyze veracity from the perspective of image manipulation.
| Fauxtography | COSMOS | MOCHEG | Post-4V | |||||||||||||||||||||
| T | F | U | O | T&F | All | T | F | U | O | T&F | All | T | F | U | O | T&F | All | T | F | U | O | T&F | All | |
| GPT-4V | 158 | 158 | 72 | 12 | 80.7 | 63.8 | 162 | 190 | 109 | 18 | 79.8 | 70.0 | 234 | 204 | 54 | 33 | 83.7 | 77.8 | 51 | 83 | 38 | 14 | 80.1 | 72.6 |
| LLaVA(7b) | 237 | 71 | 7 | 49 | 54.1 | 50.5 | 311 | 101 | 14 | 54 | 66.0 | 60.1 | 341 | 89 | 8 | 66 | 56.0 | 50.8 | 125 | 37 | 3 | 21 | 51.2 | 49.5 |
| LLaVA(13b) | 207 | 143 | 7 | 42 | 51.7 | 52.0 | 228 | 201 | 22 | 28 | 69.9 | 68.5 | 235 | 212 | 7 | 50 | 61.8 | 59.5 | 65 | 105 | 2 | 14 | 61.8 | 60.2 |
5.2.4 Taxonomy of Failure Reasons
Upon examining GPT-4V’s explanations, we identify six primary categories of failure reasons. Lack of Knowledge, especially obscure knowledge that contradicts intuition, easily leads the model to assess a statement’s truthfulness inaccurately. Similar Images refer to instances where the model associates with another image featuring a similar scene. In such cases, the model incorrectly extracts knowledge, leading to inconsistent content with the presented image. Distortion Due to Perspective and Lens, we find that GPT-4 cannot correctly understand images that are distorted due to the shooting angle or lens.. Incomplete Verification refers to the cases where the model does not fully consider the atomic claims within a multimodal claim. For example, when a claim contains both correct and incorrect subclaims, or when the textual claim is accurate but the accompanying image is misleading. Blurry or Incomplete Image can lead to an incorrect prediction due to the model’s uncertainty about what it sees. Ambiguity Claims involve semantic uncertainties, where a term could reference multiple entities. We show representative examples and explanations for each failure reason in Figure 8. Our taxonomy can lead to targeted optimization strategies. For instance, we could include more images with distortions or blurry or incomplete images from image scaling and cropping into the instruction-tuning dataset.
5.3 Ablation Studies
Impact of Images
To understand the impact of images on fact-checking, we conduct a comparative experiment where we ask GPT-4V and LLaVA (13b) to re-evaluate the veracity of textual claims without providing images. As shown in Table 3, the results are marked in red for values that exceed those in Table 2. GPT-4V experiences a decrease in accuracy across all datasets, while surprisingly, LLaVA (13b) shows higher accuracy compared to when images are included. To understand this, we further look into their explanations. As shown in Figure 4, in the 1st and 3rd cases, the model can make correct predictions based on internal knowledge without images, stating "it is unlikely that Barack Obama would be endorsing his wife’s …" and "Joe Biden is a public figure and it is highly unlikely that he would pose menacingly …" However, the model is misled when manipulated images are presented. The GPT-4V’s responses without image input are in Figure 9 in Appendix.
Reasoning First
In the experiments above, we prompt the model to first output a prediction, followed by step-by-step reasoning. We conduct a comparative experiment to assess the impact of reversing this order—starting with reasoning before prediction—on accuracy. The results are presented in Table 4, where values exceeding those in Table 2 are highlighted in red. We observe that starting with reasoning significantly increases the instances of GPT-4V expressing uncertainty. For example, the number of uncertain cases rises from 29 to 72 on Fauxtography and 10 to 38 on Post-4V. This increase directly leads to an obvious decline in Overall Accuracy, although True&False Accuracy remains relatively high. LLaVA(7b) still predicts most claims as True. LLaVA(13b) shows varying accuracy improvements across the last three datasets, though the changes are not substantial. We also note an increase in instances where the model fails to follow instructions, which may partially influence the accuracy metrics.
6 Limitations
We acknowledge the following limitations in our work. First, due to constraints in manpower and computational resources, we did not test the complete Fauxtography and COSMOS datasets. The limited number of samples may impede an accurate assessment of model characteristics. Second, we did not quantify the different capabilities within fact-checking, such as the ability of different models to recognize AI-generated images, identify faces and scenes, and retrieve contextual knowledge. Moreover, we did not investigate how fine-tuning with domain-specific data could potentially improve model performance. Future research will aim to address these gaps and provide a more comprehensive evaluation of the models.
@@ 7 Conclusion and Future Work
We investigated the capabilities of MLLMs to fact-check real-world claims, relying solely on their parametric knowledge and reasoning capabilities without external references. We proposed an evaluation framework, designed various experiments to address the research questions. Our results indicated that state-of-the-art MLLMs, such as GPT-4V, have the potential to assist professional fact-checkers: they can enhance their efficiency by providing reference predictions, valuable clues, and explanatory insights, together with confidence. However, open-source models like LLaVA fall behind. They can be misled by manipulated images and may generate hallucinations when no image is present. We empirically demonstrate that prompt ensembles (PE) and in-context learning (ICL) are able to improve model’s accuracy in detecting misinformation, while ICL is more efficient.
In future work, we plan to investigate how to enhance the fact-checking capabilities of models, focusing not only on improving accuracy but also on strengthening robustness and augmenting the knowledge reasoning and the verification abilities required for fact-checking. A potential approach is to use GPT-4V to enhance the capabilities of smaller open-source models through knowledge distillation. We also plan to explore how MLLMs can better support fact-checking when connected to external knowledge.
Appendix A More Experimental Results
Figure 9 displays additional responses, where "w. Image" indicates multimodal claims as input and "w/o. Image" denotes text-only claims as input. The last two rows show the results for LLaVA+ICL-1 and LLaVA+ICL-2, respectively. As shown in Figure 9, GPT-4V’s explanations demonstrate its awareness of missing images, employing phrases such as "there is no verified public record" or "without any provided context." This indicates that GPT-4V actively retrieves knowledge to verify the presence or context of images. Compared with no demonstrations, the results reveal that ICL enhances the model’s reasoning capabilities, incorporating more detailed analysis of images and discussions about the motivations behind their creation.
Table 5 shows the results with different prompts in prompt ensembles (PE). The accuracy of the GPT-4V model is quite consistent across various prompts. In contrast, the LLaVA model is very sensitive to the prompts, with some prompts showing significantly higher accuracy than others on the same dataset. However, no single prompt consistently performs best across all datasets. Table 6 displays the impact of different demonstrations in ICL on the LLaVA (13b) model. Demonstrations clearly enhance the model’s accuracy. However, the benefit of the second example is limited, and it does not show a significant improvement compared to the ICL-I cases.